Enter a simplified algebraic expression for the expression. sin cos (A))
Answer #1

sin (cos(4) A cos A EV => cosn = A *9 •C B A ) from, a ABC ABE Altber 81 - A2 AB2= => AB = 81- AL sinas 81-42 AB AC A sin (Co